OpenStudy (anonymous):

Verify that cos x+(pi/2)=-sin x

OpenStudy (anonymous):

this mean \[\cos(x+\frac{\pi}{2})=-\sin(x)\] right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you know the addition angle formula for cosine?

OpenStudy (anonymous):

Is that the same thing as sum and difference

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73 are you there

OpenStudy (anonymous):

yeah the "sum" formula if you like the word 'sum' instead of "addition" same thing

OpenStudy (anonymous):

Ok so i believe it's cos(A+B)=cos A cos B - sin A sin B, right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

replace \(A\) by \(x\) and \(B\) by \(\frac{\pi}{2}\) and it will drop right out

OpenStudy (anonymous):

So cos (x+pi/2)= cos x cos pi/2 - sin x sin pi/2 and then I would solve that

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you get it in one step

OpenStudy (anonymous):

Oh ok I was trying to do it a different way that's where I was confused thanks.

OpenStudy (anonymous):

yw